RESPECTIVE SPEED OF FIVE MOLECULES ARE 2, 1.5 ,1.6,1.6,AND 1.2 KM/S    PROBABLE SPEED IN KM/S    WILL BE?


Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

Probable speed =sqrt(2RT/M)
Mean speed = sqrt(8RT/pi M)
=> Probable speed /Mean Speed=sqrt(pi)/2
=>Prob .Vel =Mean Velocity *sqrt(pi)/2

You can calculate yourself now.

Thanks & Regards
Arun Kumar
Askiitians Faculty
IIT Delhi